OpenStudy (anonymous):

A student club has 11 members. How many ways can the club choose a president, vice president, and treasurer?

OpenStudy (anonymous):

Well first youchoose 3 people out of the 11. Ways = 11C3 Now the three people can be assigned diff jobs- Thats lke arranging them in 3! ways Therefor answer = 11C3 * 3! = 11!/(3!*8!) *3! = 11! / 8! = 11*10*9 = 110*9 = 990